The quality control manager of a cookie company is inspecting a batch of​ chocolate-chip cookies that has just been baked. If the production process is in​ control, the mean number of chip parts per cookie is 6.0.

How many cookies in a batch of 100 should the manager expect to discard if the company policy requires that all​ chocolate-chip cookies sold have at least four ​chocolate-chip parts?
Answer #1

Solution :

Given that ,

mean = = 6

Using Poisson probability formula,

P(X = x) = (e- * x ) / x!

P(X 4) = P(X = 4) + P(X = 5) + P(X = 6)

= (e-6 * 64) / 4! + (e-6 * 65) / 5! + (e-6 * 66) / 6!

= 0.8488

Probability = 0.8488

1000 * 0.8488 = 848.8  cookies expect